OpenStudy (anonymous):

what is 4x+4= -(x-9)

OpenStudy (anonymous):

okay. im still a little stuck

hero (hero):

In order to solve this problem, you'll need to know: 1. Distributive Rule 2. Inverse Operations 3. Basic Algebra Which are you having trouble with?

OpenStudy (anonymous):

distributive rule.

hero (hero):

1. Distributive Rule: a(b+c) = ab + ac if a = 4, b = 5, and c = 6, then: 4(5 + 6) = 4*5 + 4*6 4(11) = 20 + 24 44 = 44 This means that no matter what the values of a, b, and, c, both sides will be true. Therefore the distributive rule a(b+c) = ab + ac is true.

hero (hero):

Let me know if you understand that

hero (hero):

It also means that you can use the distributive rule to help solve algebra: x(4+5) = 4x + 5x 3(2x + 5) = 6x + 15 7(x + y) = 7x + 7y y(7x + 4) =7xy + 4y

OpenStudy (anonymous):

i got it thanks

hero (hero):

Okay, so this means you can solve the problem now.

hero (hero):

What do you get for x?

hero (hero):

By the way, also, for this, you'll need to know these rules: - x - = + - x + = - + x + = +

hero (hero):

Because if you're just distributing a negative -(2x - 4) = -2x + 4 because -1 x -4 = 4

hero (hero):

In algebra - and -1 mean the same thing

hero (hero):

so -(2x - 4) = -1(2x - 4)